# Optimal Distributed Similarity Estimation for Unitary Channels

### v1

@@ C.1.2 Technical Lemmas
For two terms defined in Eq. (77), we have the following upper bounds,
@@
where the first tensor factor corresponds to the main system $\mathcal{H}_{\rm main}$ and the second to the auxiliary system $\mathcal{H}_{\rm aux}$. Note that the vectors $|{\phi_{t,i,j}}\rangle$ and $|{\psi_{t,i,j}}\rangle$ are not required to be normalized. For convenience, let $\bm{j}\mathrel{\mathop{\ordinarycolon}}=(\bm{j}_{1},\cdots,\bm{j}_{T})$ range over all sequences of length $T$, with each $j_{t}\in{0,\ldots,d^{\prime}-1}$. Define $W_{\ell}\mathrel{\mathop{\ordinarycolon}}=\prod w_{t}$, and
@@
where $J_{H}^{(T)}$ is the Choi operator of Haar random channel defined in Eq. (18), Likewise, we have
@@
C.2 Coherent Access
We now consider the lower bound for completing DSEU with coherent access.
C.2.1 Proof of Theorem 4
We first establish some notations. Recall that any learning algorithms with coherent access can be described as follows. Consider two unknown unitary channels $U$ and $V$ acting on two separate quantum systems. On these devices, we input the initial state $|{0}\rangle\otimes|{0}\rangle$ on $\mathcal{H}^{\otimes 2}$ and interleave the channel queries with a sequence of $T$ adaptive data-processing channels $\mathcal{C}_{t}=\mathcal{C}_{t,1}\otimes\mathcal{C}_{t,2}$, $t=1,\cdots,T$. Assuming the auxiliary system is large enough, each data-processing channel can be represented as a unitary $W_{t}=W_{t,1}\otimes W_{t,2}$ [51]. After $T$ queries, the resulting states on the two devices can then be expressed as
@@ C.2.1 Proof of Theorem 4
$|{\Psi_{I,i}}\rangle=\sum_{\begin{subarray}{c}x_{1},\cdots,x_{T}\in\{0,1\}^{n}\\ y_{1},\cdots,y_{T}\in\{0,1\}^{n}\end{subarray}}\prod_{t=1}^{T}\overbrace{\left[\left(|{y_{t}}\rangle\!\langle{x_{t}}|\otimes I_{\rm aux}\right)W_{t,i}\right]|{0}\rangle}^{\text{$\mathcal{H}$}}\otimes\overbrace{|{x_{1}\cdots x_{T}}\rangle\otimes|{y_{1}\cdots y_{T}}\rangle}^{\text{$\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$}}.$ (104)
@@
where $J_{U}^{(T)}$ is the Choi operator of unitary $U^{\otimes T}$, as defined in Eq. (12), and similar to $J_{V}^{(T)}$. With the definition of Haar random channel, we have
@@
where $J_{H}^{(T)}$ is defined in Eq. (18). As we do in the proof of lower bound for incoherent access, we define a intermediate probability
@@ C.2.2 Technical Lemmas
Figure 2: The visualization of key proof steps. (a) The visualization of permutation map $\mathcal{R}$ (defined in Eq. (123)) when $T=3$. (b) The visualization of of $\mathcal{R}^{\dagger}$ when $T=3$. (c) The visualization of Eq. (126) when $T=2$. (d) The visualization of Eq. (128) when $T=2$.
@@
$\mathcal{R}\left(\Pi_{\rm sym}^{\prime(d,T)}\right)=\Pi_{\rm sym}^{(d^{2},T)},\quad\mathcal{R}\left(J_{U}^{(T)}\right)=\left[(U\otimes I)\Phi(U\otimes I)^{\dagger}\right]^{\otimes T}.$ (124)
Then, we can observe that the space of density matrices on the symmetric subspace of $(\mathcal{H}^{\otimes 2})^{\otimes T}$ can is spanned by $\alpha^{\otimes T}$, where $\alpha\in\mathcal{D}(\mathcal{H}^{\otimes 2})$ [52, 48]. Thus, to compute the lower bound, it suffices to calculate the following function

### v2

@@ B.1.2 Technical Lemmas
For the two terms defined in Eq. (35), we have the following upper bounds,
@@
where the first tensor factor corresponds to the main system $\mathcal{H}_{\rm main}$ and the second to the auxiliary system $\mathcal{H}_{\rm aux}$. Note that the vectors $|{\phi_{t,i,j}}\rangle$ and $|{\psi_{t,i,j}}\rangle$ are not required to be normalized. For convenience, let $\bm{j}\mathrel{\mathop{\ordinarycolon}}=(\bm{j}_{1},\cdots,\bm{j}_{T})$ range over all sequences of length $T$, with each $j_{t}\in\{0,\ldots,d^{\prime}-1\}$. Define $W_{\ell}\mathrel{\mathop{\ordinarycolon}}=\prod w_{t}$, and
@@
where $J_{H}^{(T)}$ is the Choi operator of Haar random channel defined in Eq. (20). Similarly, we have
@@
B.2 Coherent Access
We now consider the lower bound for completing DSEC with coherent access.
B.2.1 Proof of Theorem 2
We first establish some notations. Recall that any learning algorithms with coherent access can be described as follows. Consider two unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on two separate quantum systems. On these devices, we input the initial state $|{0}\rangle\otimes|{0}\rangle$ on $\mathcal{H}^{\otimes 2}$ and interleave the channel queries with a sequence of $T$ adaptive data-processing channels $\mathcal{C}_{t}=\mathcal{C}_{t,1}\otimes\mathcal{C}_{t,2}$, $t=1,\cdots,T$. Assuming the auxiliary system is large enough, each data-processing channel can be represented as a unitary $W_{t}=W_{t,1}\otimes W_{t,2}$ [57]. After $T$ queries, the resulting states on the two devices can then be expressed as
@@ B.2.1 Proof of Theorem 2
$|{\Psi_{I,i}}\rangle=\sum_{\begin{subarray}{c}x_{1},\cdots,x_{T}\in\{0,\cdots,d-1\}\\ y_{1},\cdots,y_{T}\in\{0,\cdots,d-1\}\end{subarray}}\prod_{t=1}^{T}\overbrace{\left[\left(|{y_{t}}\rangle\!\langle{x_{t}}|\otimes I_{\rm aux}\right)W_{t,i}\right]|{0}\rangle}^{\text{$\mathcal{H}$}}\otimes\overbrace{|{x_{1}\cdots x_{T}}\rangle\otimes|{y_{1}\cdots y_{T}}\rangle}^{\text{$\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$}}.$ (62)
@@
where $J_{U}^{(T)}$ is the Choi operator of unitary $U^{\otimes T}$, as defined in Eq. (12), and similarly for $J_{V}^{(T)}$. With the definition of Haar random channel, we have
@@
where $J_{H}^{(T)}$ is defined in Eq. (20). As we do in the proof of lower bound for incoherent access, we define an intermediate probability
@@ B.2.2 Technical Lemmas
Figure 2: The visualization of key proof steps. (a) The visualization of permutation map $\mathcal{R}$ (defined in Eq. (81)) when $T=3$. (b) The visualization of $\mathcal{R}^{\dagger}$ when $T=3$. (c) The visualization of Eq. (84) when $T=2$. (d) The visualization of Eq. (86) when $T=2$.
@@
$\mathcal{R}\left(\Pi_{\rm sym}^{\prime(d,T)}\right)=\Pi_{\rm sym}^{(d^{2},T)}.$ (82)
Then, we can observe that the space of density matrices on the symmetric subspace of $(\mathcal{H}^{\otimes 2})^{\otimes T}$ is spanned by $\alpha^{\otimes T}$, where $\alpha\in\mathcal{D}(\mathcal{H}^{\otimes 2})$ [58, 51]. Here we write $\alpha\mathrel{\mathop{\ordinarycolon}}=|{\alpha}\rangle\!\langle{\alpha}|$ for pure state $|{\alpha}\rangle$. Thus, to compute the lower bound, it suffices to calculate the following function